# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v2

@@ 5.3.2 Main Results Analysis
Table reftab:mainresults shows a comparison of the accuracy of the different methods in predicting future routes on two real-world datasets under three scenarios. Overall, our proposed RouteKG model achieves optimal accuracy on all metrics. It is observed that for all models, the route-level prediction accuracy is lower than the link-level prediction accuracy. This underscores the importance of modeling the consistency between different road links. Comparing the different models, we can find that Deep Learning-based methods achieve higher accuracy in general. This aligns with empirical findings and is further enhanced with the integration of additional information. For instance, models like NetTraj and RouteKG, which incorporate spatial data, outperform simpler models like RNN and its variants. Remarkably, RouteKG outperforms the NetTraj model and other baselines, even without extra information, which highlights the effectiveness of our approach to integrate KG for future route prediction. Comparing under different experimental settings, intuitively, introducing more Goal information progressively improves overall accuracy. In particular, RouteKG’s prediction accuracy is greatly improved after the gradual incorporation of Goal’s information, and at the same time, it also has an accuracy improvement of about 5.41% in comparison with the optimal baseline under the NoGoal condition, which proves the effectiveness of RouteKG in processing and utilizing Goal information, and its applicability under various conditions. Lastly, comparing between different datasets, the prediction accuracy on the chengdu dataset is better than that on the Shanghai dataset, likely attributed to the larger data volume of the former.
@@
To provide intuitive understanding about the prediction results, we show two qualitative example RouteKG outputs under NoGoal. Figure 7 illustrates, for each example, the last observed link alongside the estimated direction and top-3 predictions. Although certain future routes may display peculiar turns due to constraints imposed by the road network, most predicted future routes exhibit a correct heading based on the estimated goal direction. An important observation from the first example is the misalignment between the links adjacent to the last observed link and the predicted direction. Consequently, the predicted road link initially are constrained on the road network. However, as the prediction progresses, subsequent steps are adjusted to align with the route’s predicted direction.
@@
5.4 Ablation Analysis
This section analyzes the results of the ablation experiments. Specifically, we focus on the analysis performed on RouteKG under GoalD. By conducting these ablation experiments, we can gain insight into the importance of each component of the model and its contribution to the overall predictive capabilities of the model.
Figure 8 compares the performance of RouteKG with its two ablation variants. Notably, RouteKG w/o rerank removes the Rank Refinement Module. Experimental results show that removing this module significantly reduces prediction performance. This suggests the interconnected nature of link choices, emphasizing the need for a module to model route consistency and choice correlation. This highlights the module’s indispensability. Remarkably, even without reranking, RouteKG still outperforms most benchmark methods, particularly in the top 5 and top 10 predictions. This demonstrates RouteKG’s efficacy in identifying potential future routes, reinforcing the importance of integrating the ranking refinement module for enhancing top 1 predictions. In summary, while RouteKG effectively generates future route sets, incorporating a reranking module is crucial for accurately prioritizing the top 1,000 dollar predictions.
@@ 5.4 Ablation Analysis
5.5 Sensitivity Analysis
In this section, we perform a sensitivity analysis to assess the robustness and reliability of our model under various parameter settings.
@@ 5.5 Sensitivity Analysis
We investigate the model’s performance under varying lengths of future routes to be predicted, denoted as $\Gamma^{\prime}$. As depicted in Figure 9, by altering $\Gamma^{\prime}$ from 2 to 8, there is a noticeable trend of declining performance with increasing $\Gamma^{\prime}$ values. This indicates that predicting longer routes becomes progressively challenging due to an expanded candidate space and heightened uncertainty, particularly when lacking Goal information.
5.6 Efficiency Analysis
Efficient route prediction in transportation systems is paramount, necessitating prompt response for system operators and road users. To assess the model efficiency, we analyze the inference time of various models in two datasets, aiming to ascertain real-time performance capabilities. This allows us to identify the most efficient and responsive models, which is crucial for ensuring smooth user experiences and effective traffic management.
Figure 10 delineates the inference times across various models. Note that the results from Dijkstra and RCM-BC are omitted due to their overly high inference times. All baseline utilize the Spanning Route algorithm to generate future routes except the Markov model. The Markov model leverages pre-computed transition probabilities to sample and generate the top-$k$ predictions through $k$ iterations.
Figure 10: Inference time of different methods in Chengdu and Shanghai. (mean ± std)
RouteKG demonstrates remarkable efficiency, achieving average inference times of 598.01ms and 244.47ms for every 10k requests on the Chengdu and Shanghai datasets, respectively, with standard deviations of 1.21ms and 19.35ms. In contrast, the Dijkstra model, based on dynamic programming, takes over 38 seconds, and the RCM-BC model exceeds 1000 seconds, rendering them impractical for real-time systems. Models utilizing the Spanning Route algorithm (e.g., RNN, GRU, LSTM, NetTraj, RouteKG) show superior inference times, with less than 400 milliseconds in Chengdu and 250 milliseconds in Shanghai per 10k requests. RouteKG exhibits a marginally higher inference time, likely attributable to the reranking process. Nonetheless, these results highlight RouteKG’s suitability for real-time traffic applications, where rapid processing is essential for efficient transportation systems.
5.7 Case Study: Traffic Flow Estimation
In this section, we conduct a case study on traffic flow estimation to demonstrate the practical use cases of RouteKG, utilizing its potential to generate future routes for accurate traffic flow estimations. This can offer key insights into traffic pattern dynamics and enhancing the reliability of traffic flow predictions. Specifically, we adopt a sampling-based method for generating future routes to maximize the utility of the top-$k$ future routes predictions. Initially, the top-$k$ predictions are converted to a probability distribution using temperature scaling (Guo et al. 2017). Subsequently, we sample from the predicted top-$k$ future routes for each observed trajectory based on their probability distribution. The estimated link-level traffic flows are then obtained by aggregating the number of predicted future routes at the link level. To counter the effects of uncertainty, we iterate the experiments ten times and reported traffic flow estimation results in a mean±std format, focusing solely on the top-10 predictions for simplicity.
Table 4: Traffic flow estimation results on Chengdu and Shanghai dataset. (mean ± std)
@@ 5.7 Case Study: Traffic Flow Estimation
The effectiveness of traffic flow estimation with RouteKG is demonstrated using three standard regression metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the coefficient of determination ($\mathrm{R}^{2}$), as detailed in Table 4. RouteKG consistently outperforms in all metrics for both datasets, aligning with our main experiment results in Section 5.3. Notably, incorporating more Goal information leads to improved accuracy in traffic flow predictions, highlighting the strength of our approach.
@@
To summarize, these results suggest that RouteKG could also be an effective tool for traffic flow estimation, offering accurate and rapid analysis essential for real-time traffic management.
6 Conclusion
@@
Future research can extend this work in several ways. First, incorporating other spatial relations (e.g., function zones, spatial regions, etc.) with urban and road network attributes can augment the scalability and generalizability of the model. This would enable the model to provide high-performance feedback for multi-functional intelligent transportation services rapidly, adapting to different tasks promptly. Second, future work can potentially enhance the Spanning Route algorithm by integrating an n-ary tree pruning approach, offering a solution to model complexity increases exponentially with route prediction length. The optimized algorithm is anticipated to offer superior scalability and more efficient future route generation with reduced computational resources. Last but not least, future research might delve deeper into harnessing KGs for broader urban applications, such as employing KGs to integrate diverse datasets and learn interrelationships amongst them. For instance, discerning correlations between traffic patterns and population demographics might empower urban planners to envisage and anticipate the implications of varied urban development strategies.
Acknowledgment
This research is supported by the National Natural Science Foundation of China (NSFC42201502) and Seed Funding for Strategic Interdisciplinary Research Scheme at the University of Hong Kong (URC102010057).
Appendix A Notations
The adopted notations in this paper.
$\mathcal{T}$
The set of all GPS trajectories.
$G$
The road network data.
$\mathbf{G}$
The road network, represented by a Multi-Directed Graph (MultiDiGraph).
$\mathbf{V}$ / $\mathbf{E}$
The set of all nodes / edges of $\mathbf{G}$.
$v$ / $e$
An intersection (or node) / link (or edge) of a road network.
$r^{d}_{i}$
The goal direction of the $i$-th future route.
$\widehat{r^{d}_{i}}$
The estimated goal direction of the $i$-th future route.
$\mathbf{r}^{d}_{i}$
The relation embedding of the estimated goal direction of the $i$-th future route.
$e^{i}$
The $i$-th link (or edge) of a road network.
$e_{i}^{j}$
The $j$-th link (or edge) of the $i$-th route.
$e_{i}^{d,j}$
The direction of $j$-th link of the $i$-th route.
$e_{i}^{\Gamma}$
The direction of $j$-th link of the $i$-th route.
$\mathbf{e}$
The set of embeddings of all links
$\mathbf{e}^{\Gamma}_{i}$
The embedding of the last (i.e., the $\Gamma$-th) link in the $i$-th observed route.
$\mathbf{e}_{i}^{\cdot}$
The embedding of a link in the $i$-th observed route.
$\mathbf{e}_{i}^{d,\cdot}$
The direction embedding of a link in the $i$-th observed route.
$\mathbf{e}_{i,\perp}^{\Gamma}$
The hyperplane-projected embedding of the last link in the $i$-th observed route.
$\mathbf{e}_{\perp}$
The hyperplane-projected set of all candidate tail entities embeddings.
$x_{i}$
The $i$-th map-matched route on $\mathbf{G}$ of a GPS trajectory.
$x^{o}_{i}$
The $i$-th observed route on $\mathbf{G}$ of a GPS trajectory.
$x^{o,d}_{i}$
The set of links directions of the $i$-th observed route.
$x^{f}_{i}$
The $i$-th future route on $\mathbf{G}$ of a GPS trajectory.
$\widetilde{x_{i,k}^{f}}$
The $k$-th future route out of the top-$K$ generated routes of $i$-th future route.
$\mathbf{x}^{o}_{i}$ / $\widetilde{\mathbf{x}^{f}_{i}}$
The embedding of the $i$-th observed / future route.
$\mathbf{x}^{o,d}_{i}$ / $\widetilde{\mathbf{x}^{f,d}_{i}}$
The embedding of the $i$-th observed / future route’s directions.
$\widetilde{\mathbf{x}^{f}_{i,\perp^{c}}}$
The $\mathbf{p}^{c}$ hyperplane-projected embedding of the $i$-th future route.
$\widetilde{\mathbf{x}^{f}_{i,\perp^{s}}}$
The $\mathbf{p}^{s}$ hyperplane-projected embedding of the $i$-th future route.
$\{\widehat{x_{i,k}^{f}}\}_{k=1}^{K}$
The reranked $k$-th future route out of the top-$K$ generated routes of $i$-th future route.
$\mathcal{X}$
The set of all map-matched routes.
$\mathcal{X}^{o}$ / $\mathcal{X}^{f}$
The set of all observed / future routes.
$\widetilde{\mathcal{X}^{f}_{k}}$
The set of all generated $k$-th future routes out of the top-$K$ generated routes.
$\widehat{\mathcal{X}^{f}_{k}}$
The set of all reranked $k$-th future routes out of the top-$K$ reranked routes.
$\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})$
The predicted probability distribution indicating the likelihood of each link being the $\gamma$-th link of the $i$-th future route.
$\mathrm{Pr}(\widetilde{x^{f}_{i}})$
The set of predicted probability distributions for the $i$-th future route $x^{f}_{i}$.
$\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$
The set of all predicted probability distributions for future routes $\mathcal{X}^{f}$.
$\Gamma$ / $\Gamma^{\prime}$
The length of observed / future routes.
$\mathcal{F}$
The route prediction mapping function.
$K$
The number of generated future routes.
$\mathcal{M}_{d}$ / $\mathcal{M}_{kg}$ / $\mathcal{M}_{g}$ / $\mathcal{M}_{r}$
The mapping functions of the Data Preprocessing Module / Knowledge Graph Module / Route Generation Module / Rank Refinement Module.
$\Theta$
The parameter set of $\mathcal{F}$.
$\Theta_{kg}$ / $\Theta_{r}$
The parameter set of $\mathcal{M}_{kg}$ / $\mathcal{M}_{r}$.
$\mathbf{D}$
The direction label matrix.
$\mathbf{A}$
The node adjacent edges matrix.
$\mathbf{D}^{d}$
The inter-road direction matrix.
$\mathcal{G}$
The knowledge graph.
$\mathcal{E}$ / $\mathcal{R}$
The set of all entities / relations of $\mathcal{G}$.
$(h,r,t)$
A triplet within a knowledge graph $\mathcal{G}$, $h$ is the head entity, $r$ is the relation, $t$ is the tail entity.
$\mathbf{h}$ / $\mathbf{r}$ / $\mathbf{t}$
The head entity / relation / tail entity embedding.
$\mathbf{p}^{r}$
The hyperplane of the relation $r$.
$\mathbf{p}^{c}$ / $\mathbf{p}^{s}$ / $\mathbf{p}^{a}$ / $\mathbf{p}^{d}$
The hyperplane of the ConnectBy / ConsistentWith / DistanceTo / DirectionTo relation.
$\mathcal{R}^{c}$ / $\mathcal{R}^{s}$ / $\mathcal{R}^{a}$ / $\mathcal{R}^{d}$
The ConnectBy / ConsistentWith / DistanceTo / DirectionTo spatial relation.
$\delta$
The hidden dimension.
$\delta_{\mathcal{E}}$
The dimension of the entity embedding spaces.
$\delta_{\mathcal{R}^{c}}$ / $\delta_{\mathcal{R}^{s}}$ / $\delta_{\mathcal{R}^{a}}$ / $\delta_{\mathcal{R}^{d}}$
The dimension of the ConnectBy / ConsistentWith / DistanceTo / DirectionTo relation embedding spaces.
$\Delta_{\mathcal{R}^{c}}$, $\Delta_{\mathcal{R}^{c}}^{\prime}$
The set of valid (positive) triplets and the set of invalid (negative) triplets of relation $\mathcal{R}^{c}$.
$\Delta_{\mathcal{R}^{s}}$, $\Delta_{\mathcal{R}^{s}}^{\prime}$
The set of valid (positive) triplets and the set of invalid (negative) triplets of relation $\mathcal{R}^{s}$.
$\Delta_{\mathcal{R}^{a}}$, $\Delta_{\mathcal{R}^{a}}^{\prime}$
The set of valid (positive) triplets and the set of invalid (negative) triplets of relation $\mathcal{R}^{a}$.
$\Delta_{\mathcal{R}^{d}}$, $\Delta_{\mathcal{R}^{d}}^{\prime}$
The set of valid (positive) triplets and the set of invalid (negative) triplets of relation $\mathcal{R}^{d}$.
$\phi$
The scoring function of knowledge graph embedding.
$\psi$
The margin of positive and negative scores.
$\mathbf{W}_{\mathcal{E}}$
The trainable entity embedding matrix.
$\mathbf{W}_{\mathcal{R}^{c}}$ / $\mathbf{W}_{\mathcal{R}^{s}}$ / $\mathbf{W}_{\mathcal{R}^{a}}$ / $\mathbf{W}_{\mathcal{R}^{d}}$
The trainable ConnectBy / ConsistentWith / DistanceTo / DirectionTo relation embedding matrix.
$\mathbf{P}_{\mathcal{R}^{c}}$ / $\mathbf{P}_{\mathcal{R}^{s}}$ / $\mathbf{P}_{\mathcal{R}^{a}}$ / $\mathbf{P}_{\mathcal{R}^{d}}$
The trainable ConnectBy / ConsistentWith / DistanceTo / DirectionTo relation Hyperplane.
$\mathbf{r}^{f,c}_{i,margin},\mathbf{r}^{f,s}_{i,margin}$
The connection and consistent margins.
$\mathcal{B}$
The batch size.
$w_{\cdot}$
The weights for different loss functions.
$N_{A}$
The maximum number of the adjacent edges of all nodes in the $\mathbf{G}$.
$N_{d}$
The number of sections into which directions are discretized.
$\mathcal{L},\mathcal{L}_{\cdot}$
The loss functions.
Appendix B Minibatch version of the Spanning Route algorithm
@@
road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$;
NAE matrix $\mathbf{A}\in\mathbb{R}^{|\mathbf{V}|\times N_{A}}$;
the tree’s degree $n$.
Output :  Top-$K$ predicted batched future routes $\left\{\widetilde{x_{k}^{f}}\right\}_{k=1}^{K}$.
1 // Initialize the root node.
@@
21 end for
22leaves $\leftarrow$ GetLeaves(root)

### v3

@@ V-C2 Main Results Analysis
Table III shows a comparison of the accuracy of the different methods in predicting future routes on two real-world datasets under three scenarios. Overall, our proposed RouteKG model achieves optimal accuracy on all metrics. It is observed that for all models, the route-level prediction accuracy is lower than the link-level prediction accuracy. This underscores the importance of modeling the consistency between different road links. Comparing the different models, we can find that Deep Learning-based methods achieve higher accuracy in general. This aligns with empirical findings and is further enhanced with the integration of additional information. For instance, models like NetTraj and RouteKG, which incorporate spatial data, outperform simpler models like RNN and its variants. Remarkably, RouteKG outperforms the NetTraj model and other baselines, even without extra information, which highlights the effectiveness of our approach to integrate KG for future route prediction. Comparing under different experimental settings, intuitively, introducing more Goal information progressively improves overall accuracy. In particular, RouteKG’s prediction accuracy is greatly improved after the gradual incorporation of Goal’s information, and at the same time, it also has an accuracy improvement of about 5.41% in comparison with the optimal baseline under the NoGoal condition, which proves the effectiveness of RouteKG in utilizing Goal information, and its applicability under various conditions. Lastly, comparing different datasets, the prediction accuracy of the Chengdu dataset is better than that of the Shanghai dataset, likely because of the larger data size of the former.
@@
To provide an intuitive understanding of the prediction results, we show two qualitative examples of RouteKG outputs under NoGoal in Figure 7. Each contains the last observed link alongside the estimated direction and top-3 predictions. Although certain future routes may display peculiar turns due to constraints imposed by the road network, most predicted future routes exhibit a correct heading based on the estimated goal direction. An important observation from the first example is the misalignment between the links adjacent to the last observed link and the predicted direction. Consequently, the predicted road links are initially constrained on the road network. However, as the prediction progresses, subsequent steps are adjusted to align with the route’s predicted direction.
@@
V-D Ablation Analysis
This section analyzes the results of the ablation experiments. Specifically, we focus on the analysis performed on RouteKG under GoalD. By conducting these ablation experiments, we can gain insight into the importance of each component of the model and its contribution to the overall predictive capabilities.
Figure 8 compares the performance of RouteKG with its two ablation variants. Notably, RouteKG w/o rerank removes the Rank Refinement Module. Experimental results show that removing this module significantly reduces prediction performance. This suggests the interconnected nature of link choices, emphasizing the need for a module to model route consistency and choice correlation. This highlights the module’s indispensability. Remarkably, even without reranking, RouteKG still outperforms most benchmark methods, particularly in the top 5 and top 10 predictions. This demonstrates RouteKG’s efficacy in identifying potential future routes, reinforcing the importance of integrating the ranking refinement module for enhancing top-1 predictions.
@@ V-D Ablation Analysis
V-E Sensitivity Analysis
We perform a sensitivity analysis to assess the robustness and reliability of our model under various parameter settings.
@@ V-E Sensitivity Analysis
We investigate the model’s performance under varying lengths of future routes to be predicted, denoted as $\Gamma^{\prime}$. As depicted in Figure 9, by altering $\Gamma^{\prime}$ from 2 to 8, there is a noticeable trend of declining performance with increasing $\Gamma^{\prime}$ values. This indicates that predicting longer routes becomes progressively challenging due to an expanded candidate space and uncertainty, particularly when lacking Goal information.
V-F Efficiency Analysis
Efficient route prediction in transportation systems is paramount, necessitating prompt response for system operators and road users. To assess the model efficiency, we analyze the inference time of various models in two datasets, to examine model efficiencies. This allows us to identify the most efficient and responsive models, which is crucial for ensuring better user experiences and traffic management.
Figure 10 delineates the inference times across various models. Note that the results from Dijkstra and RCM-BC are omitted due to their overly high inference times. All baselines utilize the Spanning Route algorithm to generate future routes except the Markov model. The Markov model leverages pre-computed transition probabilities to sample and generate the top-$k$ predictions through $k$ iterations.
Figure 10: Inference time per 10K requests for different methods in Chengdu and Shanghai (mean ± std).
@@ V-F Efficiency Analysis
TABLE IV: Traffic flow estimation results on Chengdu and Shanghai datasets. (mean ± std)
RouteKG demonstrates remarkable efficiency, achieving average inference times of 598.01ms and 244.47ms for every 10k requests on the Chengdu and Shanghai datasets, respectively, with standard deviations of 1.21ms and 19.35ms. In contrast, the Dijkstra model, based on dynamic programming, takes over 38 seconds, and the RCM-BC model exceeds 1000 seconds, rendering them impractical for real-time systems. Models utilizing the Spanning Route algorithm (e.g., RNN, GRU, LSTM, NetTraj, RouteKG) show superior inference times, with less than 400 milliseconds in Chengdu and 250 milliseconds in Shanghai per 10k requests. RouteKG exhibits a marginally higher inference time, likely attributable to the reranking process. Nonetheless, these results highlight RouteKG’s suitability for real-time traffic applications, where rapid processing is essential.
V-G Case Study: Traffic Flow Estimation
In this section, we conduct a case study on traffic flow estimation to demonstrate the practical use cases of RouteKG, utilizing its potential to generate future routes for accurate traffic flow estimations. This can offer key insights into traffic pattern dynamics and enhance the reliability of traffic flow predictions. Specifically, we adopt a sampling-based method for generating future routes to maximize the utility of the top-$k$ future route predictions. Initially, the top-$k$ predictions are converted to a probability distribution using temperature scaling [78]. Subsequently, we sample from the predicted top-$k$ future routes for each observed trajectory based on their probability distribution. The estimated link-level traffic flows are then obtained by aggregating the number of predicted future routes at the link level. To counter the effects of uncertainty, we iterate the experiments ten times and reported traffic flow estimation results in a mean±std format, focusing solely on the top-10 predictions for simplicity.
The effectiveness of traffic flow estimation with RouteKG is demonstrated using three standard regression metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the coefficient of determination ($\mathrm{R}^{2}$), as detailed in Table IV. RouteKG consistently outperforms in all metrics for both datasets, aligning with our main experiment results in Section V-C. Notably, incorporating more Goal information leads to improved accuracy in traffic flow predictions, highlighting the strength of our approach.
@@ V-G Case Study: Traffic Flow Estimation
To summarize, these results suggest that RouteKG is effective in traffic flow estimation, offering accurate and rapid analysis essential for real-time traffic management.
VI Conclusion
@@
Future research can extend this work in several ways. First, incorporating other spatial relations (e.g., function zones, spatial regions, etc.) with urban and road network attributes can augment the scalability and generalizability of the model. This would enable the model to provide high-performance feedback for multi-functional intelligent transportation services rapidly, adapting to different tasks promptly. Second, future work can potentially enhance the Spanning Route algorithm by integrating an n-ary tree pruning approach, offering a solution to model complexity increases exponentially with route prediction length. The optimized algorithm is anticipated to offer superior scalability and more efficient future route generation with reduced computational resources. Lastly, future research could explore more deeply how KGs can be leveraged for broader urban applications, such as integrating diverse datasets and uncovering interrelationships between them. For instance, identifying correlations between traffic patterns and population demographics could enable urban planners to better anticipate the impact of different urban development strategies, ultimately fostering smarter, more sustainable cities and improving overall urban system efficiency.
Appendix
VI-A Minibatch version of the Spanning Route algorithm
@@
road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$; NAE matrix $\mathbf{A}\in\mathbb{R}^{|\mathbf{V}|\times N_{A}}$; tree’s degree $n$.
Output :  Top-$K$ predicted batched future routes $\{\widetilde{x_{k}^{f}}\}_{k=1}^{K}$.
2
3// Initialize the root node.
@@
21leaves $\leftarrow$ GetLeaves(root)